OpenStudy (anonymous):

10/5 X 12 - 17 X 15 -9=

OpenStudy (anonymous):

=(2)(12)−(17)(15)−9 =24−(17)(15)−9 =24−255−9 =−231−9 =−240

OpenStudy (anonymous):

Hope that helps!

OpenStudy (anonymous):

are you a flash fan i love that show

OpenStudy (anonymous):

@ItsKoreyyy

OpenStudy (anonymous):

I am a MAJOR flash fan!

OpenStudy (anonymous):

can you wait till the new episode of the flash i can't wait to see wally as kid flash

OpenStudy (anonymous):

@ItsKoreyyy

OpenStudy (anonymous):

:D I didn't get to watch last night episode unfortunately. I may tomorrow.

OpenStudy (anonymous):

Ask me a question and I will most likely know it, people quiz me on the flash all the time :P

OpenStudy (anonymous):

is the new firestorm black or white

OpenStudy (anonymous):

He is black :P

OpenStudy (anonymous):

who is the new villain for season 2 ?

OpenStudy (anonymous):

@ItsKoreyyy